OpenStudy (eric_d):

The coefficient of x^5 in the binomial expansion of (1+5x)^8 is the same as the coefficient of x^4 in the expansion of (a+5x)^7.Find the value of a

OpenStudy (science0229):

Let's use binomial theorem for this one. Did you learn binomial theorem, yet?

OpenStudy (science0229):

\[(a+b)^n=\left(\begin{matrix}n \\ 0\end{matrix}\right)a^n b^0+\left(\begin{matrix}n \\ 1\end{matrix}\right)a^{n-1} b^1+...+\left(\begin{matrix}n \\ n\end{matrix}\right)a^0 b^n\]

OpenStudy (science0229):

From here, we can see that the kth term of the expansion is\[\left(\begin{matrix}n \\ k-1\end{matrix}\right) a^{n-k+1} b^{k-1}\]

OpenStudy (eric_d):

ok

OpenStudy (science0229):

For the first expansion, (1+5x)^8, the term with x^5 is the fourth term.

OpenStudy (eric_d):

y

OpenStudy (science0229):

Let's substitute a=1 b=5x

OpenStudy (science0229):

and n=8

OpenStudy (science0229):

\[(1+5x)^8=\left(\begin{matrix}8 \\ 0\end{matrix}\right) (1)^8(5x)^0+...+\left(\begin{matrix}8 \\ 8\end{matrix}\right)(1)^0(5x)^8\]

OpenStudy (science0229):

From here, the kth term of this expansion is \[\left(\begin{matrix}8 \\ k-1\end{matrix}\right)(1)^{9-k}(5x)^{k-1}\]Since the term contains x^5,\[x^{k-1}=x^5\]So k=6. Therefore this is the 6th term.

OpenStudy (science0229):

And my previous answer is wrong. Sorry. My mind must've slipped or something.

OpenStudy (eric_d):

previous answer is wrong. which one

OpenStudy (science0229):

The term containing x^5 is the 6th term, not fourth term.

OpenStudy (eric_d):

OK

OpenStudy (science0229):

Now, we can find the coefficient of the 6th term.\[\left(\begin{matrix}8 \\ 6-1\end{matrix}\right)(1)^{9-6}(5x)^{5}=\left(\begin{matrix}8 \\ 5\end{matrix}\right)(5^5)x^5=(56)(5^5)x^5\]

OpenStudy (science0229):

I left it like that on purpose

OpenStudy (science0229):

You'll see why later

OpenStudy (eric_d):

what do I need to do nw

OpenStudy (science0229):

Now, we do the same thing to the second expansion, but this time, we find the term with x^4

OpenStudy (science0229):

Can you try this on your own?

OpenStudy (eric_d):

will try nw

OpenStudy (eric_d):

Just to let you know.. the given answer is 2

OpenStudy (science0229):

Ok.

OpenStudy (eric_d):

Is this correct ?